# VERA: Variational Inference Framework for Jailbreaking Large Language Models

### v1

@@ A.1 Implementation Details
Following prior work on adversarial prompting [25, 17, 3], we present the attacker system prompt used to condition the adversarial generator in Figure 4, which can be found at the end of the Appendix. This prompt guides the attacker LLM to produce input queries that elicit harmful responses from the target model.
A.2 Ablation study
We conduct an ablation study to understand the contributions of the choice of attacker LLM backbone, and the KL coefficient used in the ELBO objective. Table 5 reports the attack success rate (ASR) when using different attacker models to parameterize the prompt distribution. We observe that all three attacker LLMs—Vicuna-7B, LLaMA3-8B, and Mistral-7B—perform competitively, with Vicuna-7B achieving the highest ASR at 94.0%. These results suggest that VERA is robust to the choice of attacker architecture. The attacker LLM influences the diversity and quality of sampled prompts, and these results validate the effectiveness of prompt generation across diverse attacker families.
Table 5 presents the impact of varying the KL divergence coefficient during training. We observe that moderate KL values significantly improve Attack Success Rates, peaking at 94.0% when KL=0.8. This highlights a key trade-off: the KL term regularizes the prompt distribution to remain close to a prior, promoting diverse, fluent, and semantically meaningful prompts, rather than overfitting to narrow high-reward regions. When the KL coefficient is set to zero, the model is free to exploit only the reward signal, often leading to mode collapse or degenerate behavior. Conversely, setting the KL coefficient too high (1.2) overly constrains the prompt distribution, limiting expressivity and leading to a modest drop in performance. These results emphasize the importance of tuning the KL term to strike a balance between exploration (diversity) and exploitation (success), with KL$=0.8$ providing the best trade-off in our setting.
@@ A.2 Ablation study
A.3 Connection to Reinforcement Learning
This variational framework can also be interpreted through the lens of reinforcement learning. In VERA, the attacker LLM acts as the policy, generating prompts as actions, and the judge function serves as the reward function. Optimization is performed via policy gradient, specifically REINFORCE. Table 6 summarizes the comparison between RLHF [23] and VERA. Both frameworks involve training a model to maximize an external reward signal while staying close to a reference distribution.
Table 6: Structural comparison between RLHF and VERA.
Component RLHF VERA
Reward signal Learned $R_{\phi}(x,y)$ from human prefs Judge score $R(x)=\mathrm{J}(x,\hat{y})$
Entropy term $-\alpha\,H[\pi_{\theta}]$ $-\alpha\,H[q_{\theta}]$
Update rule PPO REINFORCE
Appendix B Experimental Design Details
@@
Appendix C Example Prompt Generations
@@ Appendix D Limitations
Appendix E Societal Impact

### v2

@@ 5 Conclusion
Acknowledgements
This research is supported in part by NSF IIS-2508145 and Amazon Research Award.
@@ A.1 Implementation Details
Following prior work on adversarial prompting [jb_prompts, liu2024autodanturbo, chao2023jailbreaking], we present the attacker system prompt used to condition the adversarial generator in Figure 4, which can be found at the end of the Appendix. This prompt guides the attacker LLM to produce input queries that elicit harmful responses from the target model.
A.2 Connection to Reinforcement Learning
This variational framework can also be interpreted through the lens of reinforcement learning. In VERA, the attacker LLM acts as the policy, generating prompts as actions, and the judge function serves as the reward function. Optimization is performed via policy gradient, specifically REINFORCE. Table 4 summarizes the comparison between RLHF [ouyang2022training] and VERA. Both frameworks involve training a model to maximize an external reward signal while staying close to a reference distribution.
Table 4: Structural comparison between RLHF and VERA.
Component RLHF VERA
[…]
Entropy term $\alpha\,H[\pi_{\theta}]$ $\alpha\,H[q_{\theta}]$
Update rule PPO REINFORCE
Appendix B Ablation Study
We conduct an ablation study to understand the effect of key components in VERA: attacker optimization, attacker model backbone, KL regularization, and the judge model used for reward feedback.
Comparison with Best-of-N
To assess whether VERA’s effectiveness stems from gradient-based optimization or from the system prompt conditioning, we compare against a Best-of-N (BoN) baseline. In this setup, we disable all parameter updates by freezing the attacker LLM and generate $N=E\times B$ prompt samples, where $E$ is the number of optimization steps and $B$ is the batch size. The highest-scoring prompt under the judge is selected as the output. The performance gap in Table 5 underscores the importance of optimization: VERA does not simply exploit system prompt priors, but learns a task-specific distribution over adversarial prompts tailored to the target behavior. Static sampling fails to achieve competitive results, highlighting that optimization enables the attacker to discover a behavior-specific prompt distribution that static sampling cannot approximate.
Table 5: Effect of VERA optimization versus Best-of-N (BoN). VERA significantly outperforms BoN, highlighting the importance of VERA optimization.
Method VERA BoN
ASR (%) 94.00 48.50
Effect of Attacker Backbone
Table 7 reports the attack success rate (ASR) when using different attacker models to parameterize the prompt distribution. We observe that all three attacker LLMs—Vicuna-7B, LLaMA3-8B, and Mistral-7B—perform competitively, with Vicuna-7B achieving the highest ASR at 94.0%. These results suggest that VERA is robust to the choice of attacker architecture. The attacker LLM influences the diversity and quality of sampled prompts, and these results validate the effectiveness of prompt generation across diverse attacker families.
Effect of KL Coefficient
Table 7 presents the impact of varying the KL divergence coefficient during training. We observe that moderate KL values significantly improve Attack Success Rates, peaking at 94.0% when KL=0.8. This highlights a key trade-off: the KL term regularizes the prompt distribution to remain close to a prior, promoting diverse, fluent, and semantically meaningful prompts, rather than overfitting to narrow high-reward regions. When the KL coefficient is set to zero, the model is free to exploit only the reward signal, often leading to mode collapse or degenerate behavior. Conversely, setting the KL coefficient too high (1.2) overly constrains the prompt distribution, limiting expressivity and leading to a modest drop in performance. These results emphasize the importance of tuning the KL term to strike a balance between exploration (diversity) and exploitation (success), with KL=0.8=0.8 providing the best trade-off in our setting.
@@ Appendix B Ablation Study
Effect of Judge Model
To assess the impact of the judge model’s calibration on VERA’s performance, we conduct an ablation study across multiple judges. Specifically, we compare:
  • HB Validation Classifier: The HarmBench validation classifier[mazeika2024harmbench].
  • Strong Reject (SR) Judge: A LLM-based classifier fine-tuned for harmfulness classification by souly2024strongreject.
  • GPT-4o Mini: A prompt-based LLM judge, where GPT-4o-mini is queried with a standardized harmfulness evaluation prompt.
Table 8 reports ASR when VERA is trained using each judge. We find that performance improves with stronger and more calibrated judges: the HB Validator achieves the highest ASR (94.0%), while the GPT-4o-based judge results in lower success due to noisier, less reliable gradients. Nevertheless, all judges support successful optimization, indicating the generality and robustness of our approach.
We also analyze whether our one-sample judge estimator may suffer from high variance or bias. Since, our estimator uses a proxy classifier to approximate harmfulness; some bias is inevitable in black-box settings, however, VERA’s consistent performance across multiple models (Table LABEL:tab:asr-main) suggests that the judge approximates the true labeling function sufficiently well in practice. For variance, we evaluate the consistency of judged harmfulness across 10 generations per prompt (Vicuna-7B, $T=0.7$) and find a low mean standard deviation (0.107), indicating stable feedback. These findings suggest that the judge-based reward signal is sufficiently reliable to guide prompt optimization.
Table 8: Effect of judge model on optimization. While stronger judges yield more reliable feedback and higher ASR, VERA consistently performs well across all judge models.
Judge GPT-4o Mini StrongReject HB Validator
ASR (%) 83.00 89.00 94.00
Appendix C Additional Results
Evaluation on AdvBench Dataset
To broaden our evaluation beyond HarmBench, we report additional results on the AdvBench dataset [zou2024improving]. Specifically, we use the 50 most harmful questions identified by chen2024when and compute ASR using a keyword-based judge as in their setup. Table 9 shows the ASR of VERA compared to existing baselines on two target models: LLaMA2-7B-Chat and Vicuna-7B. We observe that VERA outperforms all prior methods on Vicuna-7B and remains competitive on LLaMA2-7B-Chat, demonstrating its generality across different threat models and evaluation datasets.
Table 9: Attack success rate on AdvBench subset. VERA demonstrates competitive performance.
Method LLaMA2-7B-Chat Vicuna-7B
GCG 10.0 72.0
AutoDAN 12.0 82.0
GPTFuzzer 12.0 100.0
PAIR 8.0 64.0
VERA (Ours) 16.0 86.0
Evaluation on different Defenses
To further evaluate the robustness of VERA, we assess its performance against additional recently proposed defense mechanisms: LLaMA Guard [llamaguard], SmoothLLM [smoothllm], and Robustly-Aligned LLM (RA-LLM) [rallm]. We apply the attackers to the target model, and then apply the defense methods to the generated prompts and measure the by-pass rate. For each, we measure the bypass rate, i.e., the fraction of adversarial prompts that successfully evade the defense.
LLaMA Guard Evaluation
LLaMA Guard is an instruction-tuned classifier designed to flag prompts that violate behavioral constraints. We evaluate the bypass rate of adversarial prompts when passed through the guard. As shown in Table 10, VERA achieves the highest bypass rate, substantially outperforming GCG and AutoDAN. We hypothesize that template-based attacks like GCG and AutoDAN are more easily detected by LLaMA Guard, as their prompts are widely known and likely included in its training data. In contrast, VERA’s learned prompts are behavior-specific and diverse, often falling outside the guard’s training distribution, enabling greater evasiveness.
Table 10: Bypass rate (%) on LLaMA Guard. VERA significantly outperforms other attacks.
Method Bypass Rate (%)
VERA 17.50
GCG 5.06
AutoDAN 6.25
SmoothLLM Evaluation
SmoothLLM detects brittle attacks by applying random perturbations to the input and rejecting those with inconsistent classification outcomes. We evaluate robustness using the RandomSwapPerturbation scheme with 10 smoothing copies. The result in Table 11 shows AutoDAN performs best due to its handcrafted, semantically robust templates. GCG suffers from syntax fragility under perturbations. While VERA underperforms AutoDAN, it outperforms GCG due to its more grounded and diverse prompt distribution.
Table 11: SmoothLLM evaluation using randomized character swaps. VERA outperforms GCG despite lacking handcrafted prompts, indicating higher semantic robustness.
Method # Copies Perturbation Type JB Rate (%)
VERA 10 RandomSwap 58.75
GCG 10 RandomSwap 41.77
AutoDAN 10 RandomSwap 95.00
RA-LLM Evaluation
RA-LLM evaluates prompt robustness by applying token-level deletions, aiming to break adversarial intent while preserving benign content. We report bypass rates under its default threshold setting. RA-LLM’s token deletions break the fixed patterns of AutoDAN, nullifying its attack. GCG occasionally survives due to persistent suffixes. VERA also experiences degradation, but its diversity allows a small number of robust prompts to succeed.
Table 12: Bypass rate (%) under RA-LLM filtering.
Method Bypass Rate (%)
VERA 2.50
GCG 3.79
AutoDAN 0.00
Across all defenses, VERA maintains competitive or superior performance compared to both white-box (GCG) and black-box (AutoDAN) baselines. Its robustness arises from learning a semantically meaningful and diverse adversarial prompt distribution, rather than relying on brittle templates or suffixes. These results underscore the need for future defenses to account for such distributional adversarial attacks.
Appendix D Experimental Design Details
@@
Appendix E Example Prompt Generations
@@ Appendix F Limitations
Appendix G Societal Impact